Why does -10 work for 2logx=2 but not log(x^2)=2
Answer by stanbon(75887)

Why does -10 work for 2logx=2 but not log(x^2)=2
-----
2log(x) = 2
log(x) = 1
x = 10^1
x = 10
-------------------
log(x^2) = 2
2log(x) = 2
log(x) = 1
x = 10^1
x = 10
----------------------
Note:: It both cases, x = 10.
Cheers,
Stan H.
--------------